OpenStudy (anonymous):

1 =A+2B -8 =2C+B 0 =4A+C Solve for A, B & C

OpenStudy (anonymous):

Take the third equation and subtract 4A from both sides -4A = C Take -4A and substitute into the second equation so -8 = 2C + B turns into -8 = 2(-4A) + B take the first equation and the new equation in the previous line and

OpenStudy (anonymous):

... subtract the first equation from TWICE new equation -16 = -16A + 2B 1 = A + 2B to get -17 = -17A so A = 1 which means from -4A = C that -4(1) = C = -4 and from the first equation 1 = 1 + 2B or B = 0

OpenStudy (phi):

Using elimination, which is very methodical, is a good way to go 1 =A+2B -8 =2C+B 0 =4A+C use elimination. First, put the coefficients in order A B C 1 2 0 1 0 1 2 -8 4 0 1 0 multiply the first row by -4 and add to the third row 1 2 0 1 0 2 2 -8 0 -8 1 -4 multiply 2nd row by 4 and add to third row 1 2 0 1 0 2 2 -8 0 0 9 -36 backsolve (work from bottom up) C= -36/9 = -4 2B -8 = -8, B=0 A +2*0 +0* -4= 1, A=1